The volume of a cone is 3000 whath appens to the volume if u double the diameter

Answer by Theo(13342)

v1 = 1/3 * pi * r^2 * h

if you double the diameter, then you are doubling the radius as well.

v2 = 1/3 * pi * (2r)^2 * h

simplify this to get:

v2 = 1/3 * pi * 4 * r^2 * h

divide v2 by v1 to get:

v2 / v1 = (1/3 * pi * 4 * r^2 * h) / (1/3 * pi * r^2 * h)

simplify to get v2 / v1 = 4

the volume will be 4 times as large.

if v1 = 3000, then v2 = 12000.